OpenStudy (anonymous):

Evaluate the log: \[log_6 \frac{1}{36}\]

hartnn (hartnn):

can you try to write 1/36 as the exponent of 6 ?

OpenStudy (anonymous):

I can try: 6^(-2)

hartnn (hartnn):

36 = 6^2 and \(\large \dfrac{1}{a}=a^{-1}\)

hartnn (hartnn):

YES! thats correct :)

OpenStudy (anonymous):

Ok but what does it mean by evaluate? Do I write: \[log_6 \frac{1}{36}=-2\]

hartnn (hartnn):

so, \(\large \log_6 6^{-2 } = (-2) \log_66\) because of the property \(\large \log m^n = n \log m\)

hartnn (hartnn):

yes, -2 is correct!

hartnn (hartnn):

evaluate means to find the value

OpenStudy (anonymous):

Ok thank you :)

hartnn (hartnn):

welcome ^_^